# Cryoscope: A Cryogenic Infrared Survey Telescope

### v1

@@ Abstract
We present Cryoscope—a new 50 deg2 field-of-view, 1.2 m aperture, K-dark survey telescope to be located at Dome C, Antarctica. Cryoscope has an innovative optical-thermal design wherein the entire telescope is cryogenically cooled. Cryoscope also explores new detector technology to cost-effectively tile the full focal plane. Leveraging the dark Antarctic sky and minimizing telescope thermal emission, Cryoscope achieves unprecedented deep, wide, fast and red observations, matching and exceeding volumetric survey speeds from the Ultraviolet Explorer, Vera Rubin Observatory, and Nancy Grace Roman Space Telescope. By providing coverage beyond wavelengths of 2 $\,\mu$m, we aim to create the most comprehensive dynamic movie of the most obscured reaches of the Universe. Cryoscope will be a dedicated discovery engine for electromagnetic emission from coalescing compact binaries, Earth-like exoplanets orbiting cold stars, and multiple facets of time-domain, stellar and solar system science. In this paper, we describe the scientific drivers and technical innovations for this new discovery engine operating in the K-dark passband, why we choose to deploy it in Antarctica, and the status of a fifth-scale prototype designed as a Pathfinder to retire technological risks prior to full-scale implementation.
@@ 1 Introduction
Pioneering infrared surveys include IRAS (Neugebauer et al., 1984), WISE/NEOWISE (Wright et al., 2010; Mainzer et al., 2011), 2MASS (Skrutskie et al., 2006) and VISTA (Sutherland et al., 2015). Launched in the 1980s, IRAS was the first infrared survey satellite to undertake an all-sky survey. WISE/NEOWISE, launched in 2009, provided orders-of-magnitude improvements upon IRAS’ sensitivity and spatial resolution, producing the longest time-baseline catalog in the mid-infrared with repeat observations over more than a decade. In the near-infrared J (1.25 $\,\mu$m), H (1.65 $\,\mu$m), and Ks (2.16 $\,\mu$m) bandpasses that are observable from the ground, 2MASS and VISTA produced single reference surveys rather than high-cadence data. 2MASS, using two 1.3 m telescopes, operated between 1997 and 2001, generating an all-sky release of 471 million sources with high photometric accuracy. VISTA/VIRCAM, using a 4 m telescope, extended this catalog with deeper and higher-resolution observations, focusing on specific areas of the sky, and is now decommissioned.
@@
These telescopes, however, are limited in sensitivity due to their small apertures and high sky backgrounds. Space directly eliminates sky background. Euclid, launched in 2023 (Laureijs et al., 2024), and the Nancy Grace Roman Space Telescope, set to launch in 2026, will reach unmatched resolutions and depths in the infrared. However, their field-of-view and science missions do not make them suitable for dedicated, high-cadence, large-area time-domain astronomy. Roman’s 0.281 deg2 and Euclid’s 0.55 deg2 near-infrared fields-of-view, coupled with planned day-timescale cadences for some surveys (e.g., Roman’s high latitude time domain reference survey, Rose et al., 2021) serve as the primary obstacles towards increasing their volumetric survey speed.
@@
A sensitive, high-cadence and wide-field survey is essential to uncovering the dynamic infrared universe. To bridge the current capability gap, we present Cryoscope, an infrared surveyor to be operated robotically at the Concordia Station at Dome C, Antarctica. With a design capable of taking advantage of the low thermal background and seeing conditions of the site, the 1.2 m telescope will be able to perform the deepest and fastest survey ever attempted in the infrared. Cryoscope will be able to survey the entire accessible southern sky $\approx$ 20,000 deg2 to a depth of 21.9 AB mag in the K-band every 13 hours continuously in the winter.
@@
Cryoscope is designed to undertake a time-domain survey that is unparalleled in volumetric mapping speed in the infrared bands (see Figure 1). Cryoscope’s volumetric mapping speed exceeds the speed of the reddest optical filter, y-band, of the Rubin Observatory by a factor of 1.4. Cryoscope’s speed is 48×48$\times$48 × faster than reference surveys with the Roman Observatory beyond 2 $\mu$m (e.g., Galactic bulge time domain survey, Roman Core Community Survey, 2025). For the full-scale Cryoscope, the 1.2 m aperture combined with the dark Concordia sky, yields a $5\sigma$ depth of 24.1 mag AB achieved in 1 hr (its fifth-scale prototype, the Pathfinder, yields a depth of 20.2 mag in the same time). Cryoscope’s 49.6 deg2 field-of-view (see Figure 2) is 5×5$\times$5 × larger than Rubin’s and 200×200$\times$200 × larger than Roman’s. Furthermore, combining this exquisite infrared sensitivity with the continuous observing season in Antarctica enables a unique search of the southern infrared sky, facilitating a vast array of science cases from characterizing temperate and terrestrial planets to discovering explosive transient events and monitoring their light curves.
@@
Table 1: The infrared survey landscape
Band MPix Deg2 Aperture (m) Status
VISTA/VIRCAM J, H, K 64 0.6 4 Decommissioned
Palomar Gattini-IR J 4 25 0.3 Operating since 2018
WINTER Y,J,H 12.4 1.2 1.2 Operating since 2024
DREAMS J,H 6 3.9 0.5 In construction
Roman I to K 288 0.28 2.4 Launch in 2026
Cryoscope Pathfinder K 4 16 0.26 In construction
Cryoscope K 600 50 1.2 Proposed 2030
Figure 1: Volumetric survey speed of Cryoscope and the prototype Cryoscope Pathfinder, compared to the Rubin Observatory, Ultraviolet Explorer (UVEX), the Roman space telescope high latitude time domain survey (HLTDS) and Galactic bulge time domain survey (GBTDS), and Euclid space telescope. Refer to Appendix A for survey parameters in this calculation.
Figure 2: Cryoscope’s field-of-view is more than 40×40$\times$40 × greater than other surveys with coverage beyond 2 $\mu$m.
@@ 2.1 The site
The high Antarctic plateau is indispensable to observational astronomy. With a contiguous area spanning the size of Australia at elevations exceeding 2,000 m, the region is home to numerous international research stations. The South Pole station (2,835 m) currently hosts an American research program that specializes in sub-millimeter observations and high-energy particle detections. Higher locations such as Dome A (4,093 m) and Dome C (3,233 m), respectively overseen by China and France/Italy, host observatories operating at optical wavelengths. The Dome C station is particularly attractive for a survey telescope because it has people on-site year-round and is already home to robust observational programs evidenced by The Antarctic Search for Transiting Planets (ASTEP) (Guillot et al., 2015; Crouzet et al., 2020; Schmider et al., 2022; Dransfield et al., 2022). The scientific advantages of operating such facilities at this location are as follows:
  • Low sky background: As shown in Ashley et al. (1996) and Nguyen et al. (1996), the K-band sky background at the South Pole is about 40 times lower than the best temperate ground sites thanks to a combination of low temperature and low water vapor content. With the minimum at 2.4 $\mu m$ caused by the disappearance of airglow lines, the K-dark window (2.2–2.5 microns) is the one to be optimally exploited for astronomy (Figure 3; also see Li et al., 2016). At the higher elevations and colder temperatures afforded by Dome C, the sky brightness is anticipated to be even lower than at the South Pole.
@@
  • Excellent weather: The studies of Crouzet et al. (2010, 2018), which analyzed data collected over four winter seasons with the ASTEP South refractor, demonstrated the great photometric stability of the Concordia site. The analysis showed that conditions were considered photometric approximately two-thirds of the time, with uninterrupted observing periods lasting up to 100 hours during the polar winter. The effective duty cycle for observations reached up to 92%, emphasizing the station’s unparalleled ability to support continuous monitoring. The study of weather conditions at different heights at Concordia over ten years (2010-2019) by Genthon et al. (2021) shows that mean monthly temperatures on the ground are between $\text$-60 C and $\text$-65 C during the Antarctic night and temperature can reach as low as $\text$-80 C. Contrary to other locations in Antarctica and particularly on the coast, the wind conditions are very favorable, with mean monthly winds below 4.5 m/s on the ground, increasing to less than 7 m/s at 18 m elevation. These results highlight Concordia’s suitability for high-precision photometric campaigns, particularly those requiring extended and stable observing windows.
@@
[…]
@@ 2.2 Telescope design
At full scale, Cryoscope will be a 1.2 m telescope with a 50 deg2 field-of-view in the K-dark bandpass while keeping thermal background well below the very dark K-band sky in the Antarctic. This requires a radically different approach to traditional NIR imagers which block telescope radiation by re-imaging the pupil onto a cold Lyot stop or by carefully baffling room-temperature telescopes (e.g. ESO’s VISTA telescope). Both methods lead to higher background and yield much smaller fields-of-view than Cryoscope which cryogenically cools the entire optical path to reduce thermal background. The cryogenically cooled infrared detector is placed at prime focus with the evacuated telescope replacing the detector cryostat.
Figure 4: Cross-section of the telescope showing the major components of the design.
@@
Figure 5: Non-sequential ray tracing of scattered light entering the first corrector lens.
@@ 2.3 The camera
The baseline detector is HELLSTAR, HgCdTe Extremely Large Layout Sensor Technology for Astrophysics Research, a device developed in the SATIN program by Rochester Institute of Technology (RIT) and Raytheon. The detector uses HgCdTe, deposited on silicon, as the light-sensitive layer bump-bonded to a read-out integrated circuit (ROIC) originally developed by Sensor Creations, Inc., now Attollo Engineering. Figure 6 shows images of the ROIC during the functional performance testing program funded by RIT’s Center for Detectors.
Figure 6: ROIC functional performance. Left: HeNe laser light projected onto a wafer containing the ROICs. Middle: Reverse-scale image constructed from data that the central ROIC produced while illuminated by the laser light. Right: Same image but with bias subtracted and flat field correction applied. Collectively, the panels demonstrate that the ROIC is functional.
The wavelength cutoff is tuned to 2.65 $\mu$m in order to take advantage of the very low thermal background of the telescope and atmosphere. During the development projects leading to the final detector recipe, the team iterated the design across a range of pixel geometries and doping profiles (Hanold et al., 2015; Figer et al., 2018, 2022; Buntic et al., 2024). The final design delivers read noise of  7 e- in Fowler-16 (25 e- CDS), QE near 90% at the peak near 2.4 $\mu$m, and dark current of 0.04 e-/second/pixel. These performance values are all better than the requirements for Cryoscope, in part because the application is heavily background-limited. Figure 7 shows the QE, demonstrating the custom long-wavelength cutoff for Cryoscope. This device was also successfully demonstrated on-sky using the 24-inch telescope of the C.E.K. Mees Observatory. Figure 8 shows a K-band image of M42, clearly showing the Trapezium and Becklin–Neugebauer object, along with a chart showing corresponding stars from 2MASS data (Buntic et al., 2024). Table 2 gives measured performance for detectors from the SATIN program along with goals for HELLSTAR.
@@
Figure 8: Left: Image of M42 obtained with a 1Kx1K version of a SATIN detector using HgCdTe deposited on silicon. Right: A plot of object locations from the data compared to those from 2MASS from Buntic et al. (2024).
HELLSTAR has 4Kx6K pixels and could be arranged in a 9x6 mosaic of 54 detectors to create a 36Kx36K, or 1.36 gigapixel, focal plane. There are other possible detectors using Raytheon all-digital focal planes, such as the SB500 (4K×$\times$4K×$\times$18 $\,\mu$m) and SB586 (4K×$\times$4K×$\times$10 $\,\mu$m), the latter of which has been successfully acquired by the DKIST project. Teledyne and Leonardo may also bid on large format HgCdTe sensors with 10 $\,\mu$m pixels.
@@ 3.1 Description
The Cryoscope Pathfinder is an f/2, 26 cm aperture, 16 deg2 field-of-view imager which will retire technological risk for the future $1.2$-meter, 50 deg2 field-of-view design. The field-of-view imaged by the Pathfinder at 7.1 arcsec/pixel is limited by the available detector and not the optical design. The Pathfinder’s modest 260 cm aperture is selected such that its legacy H2RG detector ($2048\times 2048$ pixels, 18 $\,\mu$m/pixel) and housing obstruct less than 15% of the incident beam. The telescope’s focal ratio approximates that planned for the meter-scale Cryoscope.
@@
The Pathfinder’s principal goal is to retire the technological risk inherent to the design of an Antarctic ultra-wide-field infrared surveyor detailed in §2.2. The Pathfinder will be the first infrared surveyor deployed at Dome C to measure the K-dark sky brightness over a near-continuous winter observing season. To exploit the sensitivity gains driven by reduced sky background, the Pathfinder will also meet the requirement to minimize thermal background emission from the telescope itself to remain sky-background-limited.
@@
The bench-tested optics have been installed and aligned in the cryostat (Figure 11). The K-dark filter was manufactured by Asahi Spectra, which has $>99\%$ transmission in the passband with 50-50 points at 2.25 $\mu$m and 2.5 $\mu$m, and a center wavelength of 2.375 $\mu$m, optimized for an angle of incidence of 9.65. All cryostat elements have been baked out, and zeolite dessicants ensuring dry air space between the corrector lenses have been installed. The convex corrector lens serving as the vacuum window has been shown to hold atmospheric pressure at environmental temperatures of -$\textnormal{-}85^{\circ}$ C, with no performance loss over a 2-week test. All seals in the system have successfully held vacuum over repeated temperature cycles from ambient lab temperature to -$\textnormal{-}40^{\circ}$ C. Testing to lower temperatures is ongoing. Thermal interlocks, servos and telemetry are fully operational and have been validated during thermal cycles. The legacy H2RG detector has been installed at prime focus and is currently under test. Sunpower’s Cryotel GT Stirling cryocooler is employed to maintain the telescope and detector at operating temperatures of 100 K. Initial testing revealed higher-than-expected final detector temperatures. As the cryocooler is evaluated, the detector readout and temperature-dependent dark current properties are being tested. Planned improvements in cryocooler performance or a two-cryocooler design will reduce detector temperature and dark current to within required ranges based on estimated sky brightness levels. Higher-conductance cold thermal straps for the detector and optics core are currently being designed to further improve cryocooler performance.

### v2

@@ Abstract
We present Cryoscope—a new 50  deg2 field-of-view, 1.2  m aperture, $K_{dark}$ survey telescope to be located at Dome C, Antarctica. Cryoscope has an innovative optical-thermal design wherein the entire telescope is cryogenically cooled. Cryoscope also explores new detector technology to cost-effectively tile the full focal plane. Leveraging the dark Antarctic sky and minimizing telescope thermal emission, Cryoscope achieves unprecedented deep, wide, fast and red observations, matching and exceeding volumetric survey speeds from the Ultraviolet Explorer, Vera Rubin Observatory, Nancy Grace Roman Space Telescope, SPHEREx, and NEO Surveyor. By providing coverage beyond wavelengths of 2 $\,\mu$m, we aim to create the most comprehensive dynamic movie of the most obscured reaches of the Universe. Cryoscope will be a dedicated discovery engine for electromagnetic emission from coalescing compact binaries, Earth-like exoplanets orbiting cold stars, and multiple facets of time-domain, stellar and solar system science. In this paper, we describe the scientific drivers and technical innovations for this new discovery engine operating in the $K_{dark}$ passband, why we choose to deploy it in Antarctica, and the status of a fifth-scale prototype designed as a Pathfinder to retire technological risks prior to full-scale implementation. We plan to deploy the Cryoscope Pathfinder to Dome C in December 2026 and the full-scale telescope by 2030.
@@ 1 Introduction
Pioneering infrared surveys include IRAS (Neugebauer et al., 1984), ISO/ISOCAM (Cesarsky et al., 1996), WISE/NEOWISE (Wright et al., 2010; Mainzer et al., 2011), 2MASS (Skrutskie et al., 2006) and VISTA (Sutherland et al., 2015). Launched in the 1980s, IRAS was the first infrared survey satellite to undertake an all-sky survey, with ISO/ISOCAM performing detailed observations of selected regions in the 2.5 to 18 $\,\mu$m bandpass soon after. WISE/NEOWISE, launched in 2009, provided orders-of-magnitude improvements upon IRAS’ sensitivity and spatial resolution, producing the longest time-baseline catalog in the mid-infrared with repeat observations over more than a decade. In the near-infrared $J$ (1.25 $\,\mu$m), $H$ (1.65 $\,\mu$m), and $K_{s}$ (2.16 $\,\mu$m) bandpasses that are observable from the ground, 2MASS and VISTA produced single reference surveys rather than high-cadence data. 2MASS, using two 1.3  m telescopes, operated between 1997 and 2001, generating an all-sky release of 471 million sources with high photometric accuracy. VISTA/VIRCAM, using a 4  m telescope, extended this catalog with deeper and higher-resolution observations, focusing on specific areas of the sky, and is now decommissioned.
@@
These telescopes, however, are limited in sensitivity due to their small apertures and high sky backgrounds. Space directly eliminates sky background. Euclid (Laureijs et al., 2024), launched in 2023, and the Nancy Grace Roman Space Telescope [^1] https://roman.gsfc.nasa.gov, set to launch in 2026, will reach unmatched resolutions and depths in the infrared. However, their fields-of-view (Roman’s 0.281  deg2 to Euclid’s 0.55  deg2) and science missions do not make them suitable for dedicated high-cadence, large-area time-domain astronomy.
Wider field-of-view space-based, infrared surveyors will include SPHEREx [^2] https://spherex.caltech.edu (Doré et al., 2014) and NEO Surveyor (Mainzer et al., 2023) by the end of the decade. SPHEREx is the latest mid-scale explorer from NASA, which will deliver spectro-photometric coverage from 0.75 to 5 $\,\mu$m via linear-variable filters over a 40  deg2 total field-of-view ($3.5^{\circ}\times 3.5^{\circ}$ for each of its six detector bands). SPHEREx will not, however, undertake high-cadence exposures of the entire sky or targeted pointings upon alerts, instead delivering all-sky maps four times during its two-year nominal mission. Extending from 4-10 $\,\mu$m, NEO Surveyor with its $\sim 12$  deg2 field-of-view, set to launch by the end of the decade, will build on the legacy of WISE/NEOWISE, increasing mid-infrared field-of-view and survey speed by an order of magnitude. NEO Surveyor will repeatedly scan a region of the ecliptic plane, with ecliptic latitudes limited to ±$\pm 40$ degrees, for its science goals.
@@
A sensitive, high-cadence and wide-field survey is essential to uncovering the dynamic infrared universe. To bridge the current capability gap, we present Cryoscope, an infrared surveyor to be operated robotically at the Concordia Station at Dome C, Antarctica. With a design capable of taking advantage of the low thermal background and seeing conditions of the site, the 1.2  m telescope will be able to perform the deepest and fastest survey ever attempted in the infrared. Cryoscope will be able to survey the entire accessible southern sky $\approx$ 20,000  deg2 to a point-source depth of 21.9  mag AB in the $K$-band every 13 hours continuously in the winter.
@@
Cryoscope is designed to undertake a time-domain survey that is unparalleled in volumetric mapping speed in the infrared bands (see Figure 1). Cryoscope’s volumetric mapping speed exceeds the speed of the reddest optical filter, $y$-band, of the Rubin Observatory by a factor of 1.4. Cryoscope’s speed is 48×48$\times$48 × faster than reference surveys with the Roman Observatory (e.g., considering the reddest wavelength coverage in the Galactic bulge time domain survey [^3] https://asd.gsfc.nasa.gov/roman/comm_forum). For the full-scale Cryoscope, the 1.2  m aperture combined with the dark Concordia sky, yields a $5\sigma$ point-source depth of 24.1  mag AB achieved in 1  hr (its fifth-scale prototype, the Pathfinder, yields a depth of 20.2  mag in the same time). Cryoscope’s 49.6  deg2 field-of-view (see Figure 2) is 5×5$\times$5 × larger than Rubin’s and 200×200$\times$200 × larger than Roman’s. Furthermore, combining this exquisite infrared sensitivity with the continuous observing season in Antarctica enables a unique search of the southern infrared sky, facilitating a vast array of science cases from characterizing temperate and terrestrial planets to discovering explosive transient events and monitoring their light curves.
@@
Figure 1: Volumetric survey speed of Cryoscope and the prototype Cryoscope Pathfinder compared to the Rubin Observatory, Ultraviolet Explorer (UVEX), the Roman space telescope (Galactic bulge and high-latitude time domain surveys), Euclid space telescope, SPHEREx all-sky survey, and NEO Surveyor up to 5.2 $\,\mu$m. Refer to Appendix A for survey parameters in this calculation.
Figure 2: Cryoscope’s active field-of-view exceeds other surveys in near-infrared wavelengths between 2–2.55 $\,\mu$m. Given the linear-variable filters (LVF) on SPHEREx, two of its six detector bands provide coverage in the $K$-band (Hui et al., 2024). We approximate field-of-view in these spectral channels, excluding the slight curvature in the wavelength response across detector bands.
@@ 2.1 The site
The high Antarctic plateau is indispensable to observational astronomy. With a contiguous area spanning the size of Australia at elevations exceeding 2,000  m, the region is home to numerous international research stations. The South Pole station (2,835  m) currently hosts an American research program that specializes in sub-millimeter observations and high-energy particle detections. Higher locations such as Dome A (4,093  m) and Dome C (3,233  m), respectively overseen by China and France/Italy, host observatories operating at optical wavelengths. The Dome C station is particularly attractive for a survey telescope because it has people on-site year-round and is already home to robust observational programs including The Antarctic Search for Transiting Planets (ASTEP) (Guillot et al., 2015; Crouzet et al., 2020; Schmider et al., 2022; Dransfield et al., 2022). The scientific advantages of operating such facilities at this location are as follows:
  • Low sky background: As shown in Ashley et al. (1996) and Nguyen et al. (1996), the $K$-band sky background at the South Pole is about 40 times lower than the best temperate ground sites thanks to a combination of low temperature and low water vapor content. With the minimum at 2.4 $\,\mu$m caused by the disappearance of airglow lines, the $K_{dark}$ window (2.2–2.5 $\,\mu$m) is the one to be optimally exploited for astronomy (Figure 3; also see Li et al., 2016). At the higher elevations and colder temperatures afforded by Dome C, the sky brightness is anticipated to be even lower than at the South Pole.
@@
  • Excellent weather: The studies of Crouzet et al. (2010, 2018), which analyzed data collected over four winter seasons with the ASTEP South refractor, demonstrated the great photometric stability of the Concordia site. The analysis showed that conditions were considered photometric approximately two-thirds of the time, with uninterrupted observing periods lasting up to 100 hours during the polar winter. The effective duty cycle for observations reached up to 92%, emphasizing the station’s unparalleled ability to support continuous monitoring. The study of weather conditions at different heights at Concordia over ten years (2010-2019) by Genthon et al. (2021) shows that mean monthly temperatures on the ground are between $-60$ C and $-65$ C during the Antarctic night and temperature can reach as low as $-80$ C. Contrary to other locations in Antarctica and particularly on the coast, the wind conditions are very favorable, with mean monthly winds below 4.5 m s-1 on the ground, increasing to less than 7 m s-1 at 18  m elevation. These results highlight Concordia’s suitability for high-precision photometric campaigns, particularly those requiring extended and stable observing windows.
@@
Figure 3: Measured infrared sky background from the South Pole taken from Ashley et al. (1996) and sky transmission from Hidas et al. (2000). The $K_{dark}$ filter response was optimized to exploit the window between 2.2 and 2.5 $\,\mu$m over the Antarctic plateau, similar to the methods in Li et al. (2016). The sky brightness at Dome C is expected to be even lower due to its higher elevation and colder temperatures.
@@ 2.2 Telescope design
At full scale, Cryoscope will be a 1.2  m telescope with a 50  deg2 field-of-view in the $K_{dark}$ bandpass while keeping thermal background well below the very dark $K$-band sky in the Antarctic. This requires a radically different approach to traditional NIR imagers which block telescope radiation by re-imaging the pupil onto a cold Lyot stop or by carefully baffling room-temperature telescopes (e.g., ESO’s VISTA telescope). Both methods lead to higher background and yield much smaller fields-of-view than Cryoscope which cryogenically cools the entire optical path to reduce thermal background. The cryogenically cooled infrared detector is placed at prime focus with the evacuated telescope replacing the detector cryostat.
Figure 4: Cross-section of the Cryoscope Pathfinder, showing the major components of the design. The two corrector meniscus lenses at the entrance aperture correct for aberrations from the spherical primary mirror, with the 2nd corrector also serving as a window into an evacuated cryostat interior. Cold ellipsoidal mirror-baffles line the optics core and reflect the window’s thermal emission back to the window to limit radiative cooling. Temperature controlled external radiative baffles limit the solid angle subtended to the first optical surface and maintain the first corrector lens at ambient temperature, preventing condensation from forming on the lens surface.
@@
Figure 5: Ray tracing demonstrating Cryoscope’s baffling scheme. The second corrector element (the window) is a source of thermal radiation (green rays). The ellipsoidal mirror baffles are designed to reflect the majority of the thermal radiation (blue) back to the window which is then absorbed by the lens’ convex surface (red), preventing excessive radiative cooling.
@@ 2.3 The camera
The baseline detector is a HELLSTAR detector, HgCdTe Extremely Large Layout Sensor Technology for Astrophysics Research, a device developed in the SATIN program by Rochester Institute of Technology (RIT) and Raytheon. The detector uses HgCdTe, deposited on silicon, as the light-sensitive layer bump-bonded to a read-out integrated circuit (ROIC) originally developed by Sensor Creations, Inc., now Attollo Engineering.
The wavelength cutoff is tuned to 2.65 $\,\mu$m in order to take advantage of the very low thermal background of the telescope and atmosphere. During the development projects leading to the final detector recipe, the team iterated the design across a range of pixel geometries and doping profiles (Hanold et al., 2015; Figer et al., 2018, 2022; Buntic et al., 2024). The final design delivers read noise of  7 e- in Fowler-16 (25 e- CDS), QE near 90% at the peak near 2.4 $\,\mu$m, and dark current of 0.04 e-/second/pixel. These performance values are all better than the requirements for Cryoscope, in part because the application is heavily background-limited. Figure 6 shows the QE, demonstrating the custom long-wavelength cutoff for Cryoscope. This device was also successfully demonstrated on-sky using the 24-inch telescope of the C.E.K. Mees Observatory in Bristol, New York. Figure 7 shows a color-composite $JHK$-band image of M42, clearly showing the Trapezium and Becklin–Neugebauer object, along with a chart showing corresponding stars from 2MASS data (Buntic et al., 2024). Table 1 gives measured performance for detectors from the SATIN program along with goals for HELLSTAR.
@@
Figure 7: Left: Color-composite JHK image of M42 obtained with a 1Kx1K version of a SATIN detector using HgCdTe deposited on silicon. The central four bright stars represent the Trapezium. The red object just above the center is the Becklin-Neugebaur object, demonstrating good sensitivity in the K-band.
Right: A plot of object locations from the data compared to those from 2MASS from Buntic et al. (2024).
HELLSTAR has 4Kx6K pixels and could be arranged in a 9x6 mosaic of 54 detectors to create a 36Kx36K, or 1.36 gigapixel, focal plane. There are other possible detectors using Raytheon all-digital focal planes, such as the SB500 (4K×$\times$4K×$\times$18 $\,\mu$m) and SB586 (4K×$\times$4K×$\times$10 $\,\mu$m), the latter of which has been successfully acquired by the DKIST project.
@@ 3 Prototype development
Cryoscope’s double meniscus corrector lenses enable its infrared camera to image a wide field-of-view under one of the darkest infrared sites on Earth at Dome C. In this section, we detail the status of the technology demonstrator for this new optical design.
@@ 3.1 Description
The Cryoscope Pathfinder is an f/2, 26  cm aperture, 16  deg2 field-of-view imager which will retire technological risk for the future $1.2$-meter, 50  deg2 field-of-view design. The field-of-view imaged by the Pathfinder at 7.1  arcsec/pixel is limited by the available detector and not the optical design. The Pathfinder’s modest 26  cm aperture is selected such that its legacy H2RG detector ($2048\times 2048$ pixels, 18 $\,\mu$m/pixel) and housing obstruct less than 15% of the incident beam. The telescope’s focal ratio approximates that planned for the meter-scale Cryoscope.
@@
The Pathfinder’s principal goal is to retire the technological risk inherent to the design of an Antarctic ultra-wide-field infrared surveyor detailed in §2.2. The Pathfinder will be the first infrared surveyor deployed at Dome C to measure the $K_{dark}$ sky brightness over a near-continuous winter observing season. To exploit the sensitivity gains driven by reduced sky background, the Pathfinder will also meet the requirement to minimize thermal background emission from the telescope itself to remain sky-background-limited.
@@
The bench-tested optics have been installed and aligned in the cryostat (Figure 10). The $K_{dark}$ filter was manufactured by Asahi Spectra and has $>99\%$ transmission in the passband with 50-50 points at 2.25 $\,\mu$m and 2.5 $\,\mu$m, and a center wavelength of 2.375 $\,\mu$m, optimized for an angle of incidence of 9.65. All cryostat elements have been baked out, and zeolite dessicants ensuring dry air space between the corrector lenses have been installed. The convex corrector lens serving as the vacuum window has been shown to hold atmospheric pressure at environmental temperatures of $-85$ C, with no performance loss over a 2-week test. All seals in the system have successfully held vacuum over repeated temperature cycles from ambient lab temperature to $-40$ C. Testing to lower temperatures is ongoing. Thermal interlocks, servos and telemetry are fully operational and have been validated during thermal cycles. The legacy H2RG detector has been installed at prime focus and is currently under test. Sunpower’s Cryotel GT Stirling cryocooler is employed to maintain the telescope and detector at operating temperatures of 100 K. Initial testing revealed higher-than-expected final detector temperatures. As the cryocooler is evaluated, the detector readout and temperature-dependent dark current properties are being tested. Reducing detector temperature will reduce dark current to within required ranges based on estimated sky brightness levels. Higher-conductance cold thermal straps for the detector and optics core are currently being designed to further improve cryocooler performance. We first plan to implement a liquid nitrogen-based solution to complete lab testing and validation. We will then implement a cryocooler-based approach prior to deployment at Dome C.